OpenStudy (anonymous):

(2)(pi)(6)(6) + (2)(pi)(6)(24 =1 130.97336 is this good

OpenStudy (anonymous):

don't no if it is

OpenStudy (anonymous):

value of (pi)=3.14 put it nd solve...

OpenStudy (anonymous):

this is all the probel V = π*r^2 *h V=π*(6)^2 *24=2714.3 inches^ 3 ?? S.A.=2*π*r^2 +2*π*r*h = (2)(pi)(6)(6) + (2)(pi)(6)(24

OpenStudy (anonymous):

my Instructor just want the final answer so i did it like this (2)(pi)(6)(6) + (2)(pi)(6)(24 =1 130.97336is

OpenStudy (anonymous):

so this is good (2)(pi)(6)(6) + (2)(pi)(6)(24 =1 130.97336

OpenStudy (anonymous):

(2)(pi)(6)(6) + (2)(pi)(6)(24 =1 130.97336is correct ans. as value of (pi)=3.14 or 22/7

OpenStudy (anonymous):

ok thank you should i add that to it

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

ok